### v1

@@ 3.5 Journal Diversity Index
The diversity of journals assessed using our diversity index based on Gini impurity calculated for journals with more than 100 authors was represented in a table and a map. A summary of the diversity index for selected journals is presented in Table 6. The results in Table 6 re[resent the diversity index having values close to 1 as more diverse, and values close to 0 as less diverse.
@@
Artificial Intelligence in Medicine 0.823
Journal of Medical Internet Research 0.822
@@
JMIR Medical Education 0.808
Scientific Reports 0.795
Cureus 0.771
Journal of Medical Systems 0.765
JAMA Network Open 0.735
The Lancet Digital Health 0.716
JMIR Medical Informatics 0.707
BMC Medical Informatics and Decision Making 0.697
@@
Computers in Biology and Medicine 0.644
@@
AMIA Annual Symposium Proceedings 0.610
Journal of the American Medical Informatics Association 0.597
@@
Research Square 0.499
Journal of the American Academy of Dermatology 0.440
medRxiv 0.421
AMIA Joint Summits on Translational Science Proceedings 0.411
Journal of the American College of Radiology 0.331
Nature 0.302
Figure 4: Diversity Metric per Journal
@@ 4 Discussion
The findings of our scientometric analysis reveal significant disparities in gender and geographic representation in the research on LLMs applications to healthcare. These disparities have important implications for the development and deployment of LLMs in healthcare, particularly with regard to their generalizability, usability, and fairness.
@@ 4.1 Gender Representation
Our analysis shows that the representation of female authors in LLM research is consistently low across all datasets. Overall, females constitute only 23.71% of the authors, and their representation further diminishes in key author positions, such as first and last authors, with proportions of 20.39% and 16.60%, respectively. This underrepresentation is observed in both HICs and LMICs, though the extent of the disparity varies. In HICs, female authors make up 23.45% overall, 20.69% of first authors, and 15.62% of last authors. In LMICs, the proportions are slightly higher overall (25.13%) but lower for first authors (18.52%) and last authors (23.33%).
@@ 4.2 Geographic and Income Group Representation
The geographic distribution of LLM research contributions highlights a significant imbalance favoring authors from HICs. Our data indicate that North America (HIC) dominates the research landscape, accounting for 57.46% of the publications, followed by Europe (HIC) with 19.99%. In contrast, contributions from LMICs are markedly lower. For instance, Asia (LMIC) contributes 11.79%, and Africa (LMIC) a mere 0.29%. This skewed distribution raises concerns about the global applicability of LLMs, as models predominantly developed in and for HIC contexts may not address the unique healthcare challenges and needs of LMIC populations.
The disparity is also evident in the funding sources. The majority of research funding comes from organizations based in HICs, with North America (HIC) again leading at 60.47%, and Europe (HIC) at 18.60%. LMIC funding contributions are minimal, with Asia (LMIC) at 10.99% and Africa (LMIC) at just 0.21%. This funding imbalance can further exacerbate the inequities in research outputs, limiting the ability of LMICs to conduct and contribute to high-quality LLM research. The lack of representation from LMICs, particularly in Latin America confirming the lack of resources for research in the region [23]. These challenges represent a significant risk to the generalizability and usability of LLMs in these regions, potentially perpetuating healthcare inequities.
@@ 4.3 Journal Diversity Index
The novel journal diversity index based on Gini impurity provides a quantitative measure of diversity within scientific publishing. Our analysis reveals considerable variability in diversity across journals. Journals such as ”Artificial Intelligence in Medicine” (0.823901), ”Journal of Medical Internet Research” (0.822026), and ”PLOS Digital Health” (0.821324) exhibit high diversity, indicating a more balanced representation of authors from different genders, countries, and income groups. Conversely, journals like ”Nature” (0.302608) and ”Journal of the American College of Radiology” (0.331631) show lower diversity, highlighting the need for greater inclusivity.
@@ 4.4 Recommendations
Equitable distribution of research funds: Funding bodies and agencies should prioritize the equitable distribution of research funds, supporting projects that involve LMIC researchers and address healthcare challenges pertinent to low-resource settings. This can be achieved through targeted funding calls, quotas for LMIC participation, and the inclusion of LMIC researchers in grant review panels. Funding agencies should also consider the unique constraints and challenges faced by LMIC researchers, such as limited access to computational resources and data that are particularly common in LLM research, and provide appropriate support and accommodations[26].
@@ 4.5 Limitations
While our study provides valuable insights into the diversity of LLM research in healthcare, it is important to acknowledge several limitations. The gender of authors was inferred using the Genderize.io API based on their first names. This method may introduce biases and inaccuracies, particularly for names that are unisex, culturally specific, or less common. Additionally, it does not account for non-binary and other gender identities. Country information was collected based on authors’ affiliations, which may not accurately reflect their nationality or cultural background. This can lead to misclassification, particularly for authors with affiliations in multiple countries or those working abroad.
Our analysis was based on data extracted from PubMed [33], which primarily indexes biomedical literature. Consequently, research articles published in journals not indexed by PubMed were not included in our study. This exclusion may result in an incomplete representation of the global research landscape. The study focused on articles published between January 1, 2021, and June 16, 2024. Research trends and representation patterns may evolve over time, and future studies should consider broader temporal scopes to capture these dynamics.

### v2

@@ 3.5 Journal Diversity Index
The diversity of journals assessed using our diversity index based on Gini diversity [27] calculated for journals with more than 50 authors can be seen in table 6.
The diversity index for the selected journals reveals significant variation in the representation of different geographic regions and income groups among authors. JMIR Formative Research leads with the highest diversity index of 0.830, indicating a more balanced representation of authors across various regions and income groups. This is closely followed by the Journal of Medical Internet Research and PLOS Digital Health, both of which also show high diversity scores.
On the other end of the spectrum, prestigious journals like Nature and Journal of the American College of Radiology have lower diversity indices, 0.315 and 0.356 respectively, suggesting a more concentrated representation from specific regions or groups, which might indicate a potential bias in the geographical distribution of research contributions.
@@
JMIR Formative Research 0.830
Journal of Medical Internet Research 0.825
@@
Artificial Intelligence in Medicine 0.818
JMIR Medical Education 0.809
Scientific Reports 0.804
Cureus 0.785
Journal of Medical Systems 0.766
International Journal of Medical Informatics 0.760
The Lancet Digital Health 0.760
JMIR Medical Informatics 0.742
BMC Medical Informatics and Decision Making 0.717
JAMA Network Open 0.713
@@
JAMIA Open 0.683
Studies in Health Technology and Informatics 0.673
Computers in Biology and Medicine 0.645
@@
AMIA Annual Symposium Proceedings 0.611
Journal of the American Medical Informatics Association 0.593
@@
arXiv 0.497
Research Square 0.487
Journal of the American Academy of Dermatology 0.448
medRxiv 0.429
AMIA Joint Summits on Translational Science Proceedings 0.412
Journal of Experimental Orthopaedics 0.407
Journal of the American College of Radiology 0.356
Nature 0.315
@@ 4 Discussion
The findings of our scientometric analysis reveal disparities in gender and geographic representation in the research on LLMs applications to healthcare. These disparities have important implications for the development and deployment of LLMs in healthcare, particularly with regard to their generalizability, usability, and fairness.
@@ 4.1 Gender Representation
Our analysis shows that the representation of female authors in LLM research is consistently low across all datasets. Overall, female authors constitute only 23.84% of the authors, and their representation further diminishes in key author positions, such as first being 3.06% compared with 11.9% male authors, and last authors being 2.71% compared with 14.04% male authors. This under-representation is observed in both HICs and LMICs. In HICs, female authors make up 23.55% overall, being 2.99% first authors, and 2.56% last authors. In LMICs, the proportions are slightly higher, being 25.27% overall, 3.57% first authors, and 3.66% last authors.
@@ 4.2 Geographic and Income Group Representation
The geographic distribution of LLM research contributions highlights a significant imbalance favoring authors from HICs. Our data indicate that North America (HIC) dominates the research landscape, accounting for 57.07% of the publications, followed by Europe (HIC) with 20.17%. In contrast, contributions from LMICs are markedly lower. For instance, Asia (LMIC) contributes 11.90%, and Africa (LMIC) 0.31%. This skewed distribution raises concerns about the global applicability of LLMs, as models predominantly developed in and for HIC contexts may not address the unique healthcare challenges and needs of LMIC populations.
The disparity is also evident in the funding sources. Unsurprisingly, the majority of research funding comes from organizations based in HICs, with North America (HIC) again leading at 59.15%, and Europe (HIC) at 20.42%. LMIC funding contributions are minimal, with Asia (LMIC) at 11.18%, South America (LMIC) 0.39% and Africa (LMIC) at just 0.19%. This funding imbalance can further exacerbate the inequities in research outputs, limiting the ability of LMICs to conduct and contribute to high-quality LLM research. The lack of representation from LMICs, particularly in Latin America, likely arise from the lack of resources for research in the region [28]. These challenges represent a significant risk to the generalizability and usability of LLMs in these regions, potentially perpetuating healthcare inequities.
To address these disparities, it is imperative that HICs invest more substantially in funding research in LMICs. Investment from HICs into LMIC research is not only a matter of equity but also of mutual benefit. By funding research in LMICs, HICs can gain access to unique insights and findings that are directly relevant to the global population, including in areas such as computational efficiency, data availability and diversity, and applications such as health practices that may be underrepresented in HIC-centric research.
Moreover, the lack of diverse research inputs from LMICs diminishes the potential impact and innovation that could be achieved if these perspectives were included. Investing in LMIC research will not only strengthen the global health research landscape but will also enhance the generalizability and applicability of research findings across different socio-economic and cultural contexts. In the long run, such investments will contribute to more inclusive and effective LLMs, which can better serve the global community, including populations in HICs that are increasingly diverse and interconnected with global health challenges.
@@ 4.3 Journal Diversity Index
The novel journal diversity index based on Gini diversity [27] provides a quantitative measure of diversity within scientific publishing. Our analysis reveals considerable variability in diversity across journals. Journals such as ”JMIR Formative Research” (0.830), ”Journal of Medical Internet Research” (0.825), and ”PLOS Digital Health” (0.821) exhibit high diversity, indicating a more balanced representation of authors from different genders, countries, and income groups. Conversely, journals like ”Nature” (0.315) and ”Journal of the American College of Radiology” (0.356) show lower diversity, highlighting the need for greater inclusivity.
One possible hypothesis for these disparities in diversity across journals may relate to the internal policies and practices of the journals themselves. Journals with higher diversity indices may have more explicit and proactive policies promoting inclusivity, such as encouraging submissions from underrepresented groups, including authors from LMICs, and adopting editorial practices that prioritize a diverse range of perspectives. These journals might also have more diverse editorial boards, which can influence the types of research that are accepted and the diversity of the authorship. For example, journals like ”JMIR Formative Research” and ”PLOS Digital Health” may actively work to create an inclusive environment by lowering barriers to entry for authors from diverse backgrounds, including offering fee waivers.
@@ 4.4 Recommendations
Equitable distribution of research funds: Funding bodies and agencies should prioritize the equitable distribution of research funds, supporting projects that involve LMIC researchers and address healthcare challenges pertinent to low-resource settings. This can be achieved through targeted funding calls for LMIC participation, and the inclusion of LMIC researchers in grant review panels. Funding agencies should also consider the unique constraints and challenges faced by LMIC researchers, such as limited access to computational resources and data that are particularly common in LLM research, and provide appropriate support and accommodations[31].
@@ 4.5 Limitations
While our study provides valuable insights into the diversity of LLM research in healthcare, it is important to acknowledge several limitations. The gender of authors was inferred using the Genderize.io API based on their names’ information. This method may introduce biases and inaccuracies, particularly for names that are unisex, culturally specific, or less common. Additionally, it does not account for non-binary and other gender identities. Country information was collected based on authors’ affiliations, which may not accurately reflect their nationality or cultural background. This can lead to misclassification, particularly for authors with affiliations in multiple countries or those working abroad.
Our analysis was based on data extracted from PubMed [38], which primarily indexes biomedical literature, but also include sources such as ArXiv or MedXiv. Consequently, research articles published in journals not indexed by PubMed or computational conference papers were not included in our study. This exclusion may result in an incomplete representation of the global research landscape. The study focused on articles published between January 1, 2021, and July 1, 2024. Research trends and representation patterns may evolve over time, and future studies should consider broader temporal scopes to capture these dynamics.